# EDINET-Bench: Evaluating LLMs on Complex Financial Tasks using Japanese Financial Statements

### v1

@@ Acknowledgments
We would like to thank Shimpei Fukagai, Shengran Hu, and Robert Tjarko Lange for helpful discussions.
@@ Appendix A Instruction prompt for each task
Figures 7, 8, 9, and 10 show the prompts used in our experiments.
Prompt for accounting fraud detection Please analyze the following information extracted from a Japanese company’s securities report for any signs of fraudulent activities. Please note that some data may be missing and represented as "-" due to parsing errors. The report has been verified by a certified public accountant, and the numerical values are consistent and correct from a calculation perspective. Therefore, please focus your analysis on non-numerical inconsistencies or logical red flags that could suggest fraud. Respond in the following format: JSON: “‘json { "reasoning": "string", "prob": float (between 0 and 1, probability that the report is fraudulent), "prediction": int (0: No fraud, 1: fraud) } “‘ The extracted securities report is as follows:
Figure 7: Prompt for accounting fraud detection.
Prompt for earnings forecasting Please predict whether the "親会社株主に帰属する当期純利益" (Net income attributable to owners of the parent) in the next fiscal year’s securities report will increase compared to the current fiscal year, based on the information available in the current year’s securities report. - The input is extracted from a Japanese company’s securities report. - Some information may be missing and represented as "-" due to parsing errors. - Some attributes are missing and the total does not equal the sum of the parts. Respond in the following format: JSON: “‘json { "reasoning": "string", "prob": float (between 0 and 1, probability that the profit will increase), "prediction": int (0: Decrease, 1: Increase) } “‘ The current year’s extracted securities report is as follows:
@@
Prompt for industry prediciton Based on the following financial report, classify the company into one of the Japanese industry categories. You must classify the company into exactly one of these categories: ["食品", "電気・ガス・エネルギー資源", "建設・資材", "素材・化学", "医薬品", "自動車・輸送機", "鉄鋼・非鉄", "機械", "電機・精密", "情報通信・サービスその他", "運輸・物流", "商社・卸売", "小売", "銀行", "金融(除く銀行)", "不動産"] - The input is extracted from a Japanese company’s securities report. - Some information may be missing due to parsing errors. Respond in the following format: JSON: “‘json { "reasoning": str (reasoning for the classification), "prediction": str (predicted industry category in Japanese) } “‘ The current year’s extracted securities report is as follows:
@@
Prompt for company name prediciton Please predict the name of company of the securities report, based on the information available in the current year’s securities report. - The input is extracted from a Japanese company’s securities report. - Some information may be missing and represented as "-" due to parsing errors. - Some attributes are missing and the total does not equal the sum of the parts. Respond in the following format: JSON: “‘json { "reasoning": "string", "prediction": "string" } “‘ The current year’s extracted securities report is as follows:
@@
Appendix B Example of parsed annual reports
The results of extracting BS, CF, PL, Summary, and Text items from the annual report (EDINET ID: E39409, DOC ID: S100VGBV) using edinet2dataset are shown in Figure 11, 12, 13, 14, and 15.
BS {"現金及び預金": {"CurrentYear": "1935112000"}, "現金及び現金同等物": {"Prior1Year": "1009777000", "CurrentYear": "1871411000"}, "売掛金": {"CurrentYear": "273866000"}, "仕掛品": {"CurrentYear": "5321000"}, "その他": {"CurrentYear": "36103000"}, "貸倒引当金": {"CurrentYear": "-205000"}, "流動資産": {"CurrentYear": "2240284000"}, "建物及び構築物(純額)": {"CurrentYear": "56996000"}, "機械装置及び運搬具(純額)": {"CurrentYear": "5487000"}, "土地": {"CurrentYear": "25695000"}, "減価償却累計額": {"CurrentYear": "-28267000"}, "有形固定資産": {"CurrentYear": "103401000"}, "無形固定資産": {"CurrentYear": "100789000"}, "投資有価証券": {"CurrentYear": "154996000"}, "その他、投資その他の資産": {"CurrentYear": "1784000"}, "投資その他の資産": {"CurrentYear": "263960000"}, "固定資産": {"CurrentYear": "468151000"}, "総資産": {"CurrentYear": "2708435000"}, "買掛金": {"CurrentYear": "18385000"}, "短期借入金": {"CurrentYear": "60000000"}, "1年内返済予定の長期借入金": {"CurrentYear": "263334000"}, "未払法人税等": {"CurrentYear": "33519000"}, "流動負債": {"CurrentYear": "545541000"}, "長期借入金": {"CurrentYear": "435074000"}, "非流動負債": {"CurrentYear": "453872000"}, "負債": {"CurrentYear": "999413000"}, "資本金": {"CurrentYear": "629867000"}, "資本剰余金": {"CurrentYear": "2003537000"}, "利益剰余金": {"CurrentYear": "-924382000"}, "株主資本": {"CurrentYear": "1709021000"}, "純資産": {"Prior1Year": "636232000", "CurrentYear": "1709021000"}, "負債及び純資産": {"CurrentYear": "2708435000"}}
Figure 11: A balance sheet extracted from a securities report using edinet2dataset.
CF {"当期利益": {"CurrentYear": "13054000"}, "税引前当期純利益": {"CurrentYear": "18083000"}, "減価償却費及び償却費": {"CurrentYear": "25106000"}, "貸倒引当金の増減額(△は減少)": {"CurrentYear": "23000"}, "受取利息及び受取配当金": {"CurrentYear": "-156000"}, "支払利息": {"CurrentYear": "7331000"}, "売上債権の増減額(△は増加)": {"CurrentYear": "-42526000"}, "棚卸資産の増減額(△は増加)": {"CurrentYear": "-2231000"}, "仕入債務の増減額(△は減少)": {"CurrentYear": "5663000"}, "その他": {"CurrentYear": "-1570000"}, "利息及び配当金の受取額": {"CurrentYear": "156000"}, "利息の支払額": {"CurrentYear": "-7331000"}, "営業キャッシュフロー": {"CurrentYear": "27817000"}, "投資有価証券の取得による支出": {"CurrentYear": "-129996000"}, "投資キャッシュフロー": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "-378161000"}, "長期借入れによる収入": {"CurrentYear": "190000000"}, "長期借入金の返済による支出": {"CurrentYear": "-14812000"}, "財務キャッシュフロー": {"CurrentYear": "1211978000"}, "現金及び現金同等物の増減額": {"CurrentYear": "861634000"}, "現金及び現金同等物": {"Prior1Year": "1009777000", "CurrentYear": "1871411000"}}
Figure 12: A cash flow statement extracted from a securities report using edinet2dataset.
PL {"売上高": {"CurrentYear": "1594038000"}, "売上原価": {"CurrentYear": "562526000"}, "売上総利益又は売上総損失(△)": {"CurrentYear": "1031512000"}, "販売費及び一般管理費": {"CurrentYear": "984255000"}, "営業利益": {"CurrentYear": "47256000"}, "受取利息": {"CurrentYear": "156000"}, "その他": {"CurrentYear": "0"}, "営業外収益": {"CurrentYear": "1671000"}, "支払利息": {"CurrentYear": "7331000"}, "営業外費用": {"CurrentYear": "30274000"}, "経常利益": {"CurrentYear": "18653000"}, "特別損失合計": {"CurrentYear": "570000"}, "税引前利益": {"CurrentYear": "18083000"}, "法人税、住民税及び事業税": {"CurrentYear": "7350000"}, "法人税等調整額": {"CurrentYear": "-2322000"}, "法人所得税費用": {"CurrentYear": "5028000"}, "当期利益": {"CurrentYear": "13054000"}, "親会社株主に帰属する当期純利益": {"CurrentYear": "13054000"}}
Figure 13: A profit and loss statement extracted from a securities report using edinet2dataset.
Summary {"売上高": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "1594038000"}, "経常利益": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "18653000"}, "親会社株主に帰属する当期純利益": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "13054000"}, "包括利益": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "13054000"}, "純資産額": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "1709021000"}, "総資産額": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "2708435000"}, "1株当たり純資産額": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "333.48"}, "1株当たり当期純利益又は当期純損失": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "2.67"}, "潜在株式調整後1株当たり当期純利益": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "-"}, "自己資本比率": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "0.631"}, "自己資本利益率、経営指標等": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "0.008"}, "株価収益率": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "244.6"}, "営業活動によるキャッシュ・フロー": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "27817000"}, "投資活動によるキャッシュ・フロー": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "-378161000"}, "財務活動によるキャッシュ・フロー": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "1211978000"}, "現金及び現金同等物の残高": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "1871411000"}, "従業員数": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "121"}, "平均臨時雇用人員": {"Prior4Year": "-", "Prior3Year": "-", "Prior2Year": "-", "Prior1Year": "-", "CurrentYear": "4"}}
Figure 14: A summary extracted from a securities report using edinet2dataset.
Text {"沿革": {"FilingDate": " 2 【沿革】 年月概要2010年9月静岡県焼津市にて資本金5,000千円で当社設立2011年4月感情解析エンジン「Social Emotion Engine」を開発同エンジン搭載のTwitterアプリ「Feel on!」を提供開始2012年2月東京都江東区青海に本社を移転2014年5月東京都千代田区外神田へ本社を移転2014年10月ビジネスチャットサービス「direct」を提供開始2014年12月チャットボット開発環境「daab SDK」を公開2015年10月「direct」に顧客が協力会社と安全につながるオプションサービス「direct Guest Mode(ダイレクトゲストモード)」を提供開始2016年3月徳島県徳島市に徳島ラボを開設2016年10月働き方改革支援ソリューション「direct Smart Working Solution」を提供開始2016年12月情報セキュリティマネジメントシステム(ISMS)のISO/IEC27001及びISO/IEC27017認証を取得2017年1月チャットボットレンタルサービス「direct bot RENTAL」を提供開始2018年5月大阪府大阪市西区に関西支社を開設2018年7月ユーザーの思考に合わせて進化するFAQソリューション「AI-FAQボット」を提供開始2019年5月東京都千代田区岩本町へ本社を移転2019年6月第三者割当増資により総額2.8億円を資金調達2019年10月福岡県福岡市中央区に九州支社を開設2021年5月第三者割当増資により総額約10億円を資金調達2021年10月関西支社を大阪府大阪市西区に移転2022年4月タスク管理、スケジュール管理、掲示板を搭載した「direct Apps」を提供開始2023年3月「direct Apps」に日程調整アプリ「トリスケ」を追加で提供開始2023年6月「タグショット/タグアルバム」を提供開始2024年3月東京証券取引所グロース市場に株式を上場及び公募増資(資本金537,184千円)2024年4月オーバーアロットメントによる売出しに係る第三者割当増資(資本金629,867千円)2024年11月株式会社システム・エムズの株式の取得(子会社化) "}, "事業の内容": {"FilingDate": " 3 【事業の内容】当社グループは、当社及び連結子会社1社(株式会社システム・エムズ)の計2社で構成しており、DXソリューション事業を展開しております。 当社グループは、「アイデアとテクノロジーで人々を笑顔にする!」をミッションに掲げ、「ココロ踊るチャレンジと心からの感謝にあふれる会社であり続ける」をビジョンとしております。また、役職員向けの行動指針(バリュー)として、「顧客志向」「チャレンジ」「スピード」「チームワーク」「尊重・信頼」「プロフェッショナル」の6つを定義し、徹底した顧客志向のもと、業務課題を深く理解し、課題解決を実現するためにデジタルサービスの開発・提供をしております。社名である「L is B(エルイズビー)」は、働く人々の人生の彩り・潤いになるサービスを提供したいという想いから、Life is Beautifulの頭文字を由来としてい […] 連結子会社の数及び主要な連結子会社の名称": {"CurrentYear": "(1) 連結子会社の数 1社連結子会社の名称株式会社システム・エムズ株式会社システム・エムズについては、当連結会計年度において新たに株式を取得したため、連結の範囲に含めております。" }}
Figure 15: Text items extracted from a securities report using edinet2dataset, partially omitted due to space constraints.
Appendix C Accounting fraud detection
@@
Figure 16 shows the prompt used to assess whether amendment reasons are related to accounting fraud violations. Figure 17 and 18 shows the confusion matrices for each model. Figure 19 and 20 show examples of model responses.
Prompt to detect if the reason of the amendment is related to accounting fraud 以下のテキストは訂正有価証券報告書の冒頭部分です。この訂正有価証券報告書が不適切会計、粉飾決算、会計不正に関連しているかどうかを判断してください。 特に「提出理由」の部分に着目し、以下のような言葉や表現がある場合は不正会計の可能性が高いと考えられます: - 不適切会計 - 会計不正 - 不正行為 - 粉飾決算 - 会計処理の誤り - 売上の過大計上 - 費用の過少計上 - 資産の過大評価 - 不適切な収益認識 - 監査法人からの指摘 - 社内調査 - 第三者委員会 訂正の理由が単純な記載ミスや軽微な修正ではなく、会計上の重大な問題を示している場合は「Yes」と回答し、詳細な説明を提供してください。 特に財務諸表(貸借対照表、損益計算書、キャッシュフロー計算書など)の数値に変更が生じた事例に注目してください。 会計不正とは関係ない場合や財務諸表に変更が生じていない場合は「No」と回答し、その理由を簡潔に説明してください。 以下のJSON形式で回答してください。このJSONは必ず有効なJSON形式である必要があります: “‘json { "is_accounting_fraud": bool # true or false "explanation": str # 理由を説明する "company_name": str # 会社名 } “‘ 回答は必ずこの形式に一致させてください。 JSONは必ず上記の形式の“‘json“‘タグで囲んで提供してください。 訂正有価証券報告書のテキスト: (The following text is the beginning of a corrected securities report. Please determine whether this corrected securities report is related to inappropriate accounting, window dressing (fraudulent accounting), or accounting fraud. In particular, focus on the “Reason for Submission” section. If it contains any of the following words or expressions, it is highly likely to indicate fraudulent accounting: - Inappropriate accounting - Accounting fraud - Fraudulent acts - Window dressing - Errors in accounting treatment - Overstatement of sales - Understatement of expenses - Overvaluation of assets - Improper revenue recognition - Indications from the audit firm - Internal investigation - Third-party committee If the reason for the correction is not a simple clerical mistake or minor revision but indicates a serious accounting issue, respond with “Yes” and provide a detailed explanation. Pay special attention to cases where figures in financial statements (such as the balance sheet, income statement, or cash flow statement) have changed. If it is unrelated to accounting fraud or there are no changes to the financial statements, respond with “No” and briefly explain the reason. Please respond in the following JSON format. This JSON must be valid: “‘json { "is_accounting_fraud": bool # true or false "explanation": str # Provide an explanation of the reason "company_name": str # Company name } “‘ Make sure your response matches this exact format. Wrap your JSON response with the json tag as shown above. Text of the corrected securities report:)
@@
(a) Summary
(b) BS+CF+PL+Summary
(c) BS+CF+PL+Summary+Text
(d) Summary
(e) BS+CF+PL+Summary
(f) BS+CF+PL+Summary+Text
(g) Summary
(h) BS+CF+PL+Summary
(i) BS+CF+PL+Summary+Text
(j) Summary
(k) BS+CF+PL+Summary
(l) BS+CF+PL+Summary+Text
(m) Summary
(n) BS+CF+PL+Summary
(o) BS+CF+PL+Summary+Text
Figure 17: Confusion matrix on accounting fraud detection with Anthropic and DeepSeek models.
(a) Summary
(b) BS+CF+PL+Summary
(c) BS+CF+PL+Summary+Text
(d) Summary
(e) BS+CF+PL+Summary
(f) BS+CF+PL+Summary+Text
(g) Summary
Figure 18: Confusion matrix on accounting fraud detection with OpenAI and logistic models.
@@
Figure 19: Accounting fraud detection analysis using Claude 3.7 Sonnet based on extracted data (BS, CF, PL, Summary, Text) from the securities report (EDINET Code: E00032, DOC ID: S100LQX5). Ground truth: Non-fraud.
@@
Figure 20: Accounting fraud detection analysis using GPT-4o based on extracted data (BS, CF, PL, Summary, Text) from the securities report (EDINET Code: E00032, DOC ID: S100LQX5). Ground truth: Non-fraud.
Appendix D Earnings forecasting
Figure 21 and 22 shows the confusion matrices for each model. Figure 23 and 24 shows examples of model response.
(a) Summary
(b) BS+CF+PL+Summary
(c) BS+CF+PL+Summary+Text
(d) Summary
(e) BS+CF+PL+Summary
(f) BS+CF+PL+Summary+Text
(g) Summary
(h) BS+CF+PL+Summary
(i) BS+CF+PL+Summary+Text
(j) Summary
(k) BS+CF+PL+Summary
(l) BS+CF+PL+Summary+Text
(m) Summary
(n) BS+CF+PL+Summary
(o) BS+CF+PL+Summary+Text
Figure 21: Confusion matrix on earnings forecasting with Anthropic and DeepSeek models.
(a) Summary
(b) BS+CF+PL+Summary
(c) BS+CF+PL+Summary+Text
(d) Summary
(e) BS+CF+PL+Summary
(f) BS+CF+PL+Summary+Text
(g) Summary
(h) Naive prediction
Figure 22: Confusion matrix on earnings forecasting with OpenAI, logistic, and naive prediction models.
@@ Appendix D Earnings forecasting
Figure 23: Earnings forecasting analysis using Claude 3.7 Sonnet based on extracted data (BS, CF, PL, Summary, Text) from the securities report (EDINET Code: E35466, Doc ID: S100LP1H). Ground truth: Increase.
@@
Figure 24: Earnings forecasting analysis using GPT-4o based on extracted data (BS, CF, PL, Summary, Text) from the securities report (EDINET Code: E35466, Doc ID: S100LP1H). Ground truth: Increase.
Appendix E Industry prediction
Table 9 shows the list of the 33 and 17 class labels we use for industry prediction. Note that as we explained in Section 3.3, we further merged “Energy Resources” and “Electricity & Gas” into a single class in our benchmark due to severe class imbalance.
One dataset construction issue is that SICC will reassign the industry classification when a company’s business focus shifts significantly, e.g., when the sales in the new sector increases or when mergers or business transfers cause significant changes. Since EDINET only provides the latest industry classification, we may potentially have label noise when using previous reports. Our implementation excludes companies that changed their industry label in the past based on information from the Japan Exchange Group [^11] https://www.jpx.co.jp/sicc/sectors/index.html.
Figures 25 shows the confusion matrices. Figure 26 and 27 show examples of model response.
Table 9: Mapping of SICC 33-sector classification to the TOPIX-17 Series. Note that we further merged “Energy Resources” and “Electricity & Gas” into a single class in our benchmark due to severe class imbalance.

### v2

@@ Acknowledgments
We would like to thank Ryuichi Maruyama, Shimpei Fukagai, Shengran Hu, and Robert Tjarko Lange for helpful discussions.
@@ Appendix A Instruction prompt for each task
Figures 3, 5, and 6 show the prompts used in the evaluation of the respective tasks.
Prompt for earnings forecasting Please predict whether the ”親会社株主に帰属する当期純利益” (Net income attributable to owners of the parent) in the next fiscal year’s securities report will increase compared to the current fiscal year, based on the information available in the current year’s securities report. - The input is extracted from a Japanese company’s securities report. - Some information may be missing and represented as ”-” due to parsing errors. - Some attributes are missing and the total does not equal the sum of the parts. Respond in the following format: JSON: “‘json { ”reasoning”: ”string”, ”prob”: float (between 0 and 1, probability that the profit will increase), ”prediction”: int (0: Decrease, 1: Increase) } “‘ The current year’s extracted securities report is as follows:
@@
Prompt for industry prediciton Based on the following financial report, classify the company into one of the Japanese industry categories. You must classify the company into exactly one of these categories: [”食品”, ”電気・ガス・エネルギー資源”, ”建設・資材”, ”素材・化学”, ”医薬品”, ”自動車・輸送機”, ”鉄鋼・非鉄”, ”機械”, ”電機・精密”, ”情報通信・サービスその他”, ”運輸・物流”, ”商社・卸売”, ”小売”, ”銀行”, ”金融(除く銀行)”, ”不動産”] - The input is extracted from a Japanese company’s securities report. - Some information may be missing due to parsing errors. Respond in the following format: JSON: “‘json { ”reasoning”: str (reasoning for the classification), ”prediction”: str (predicted industry category in Japanese) } “‘ The current year’s extracted securities report is as follows:
@@
Prompt for company name prediciton Please predict the name of company of the securities report, based on the information available in the current year’s securities report. - The input is extracted from a Japanese company’s securities report. - Some information may be missing and represented as ”-” due to parsing errors. - Some attributes are missing and the total does not equal the sum of the parts. Respond in the following format: JSON: “‘json { ”reasoning”: ”string”, ”prediction”: ”string” } “‘ The current year’s extracted securities report is as follows:
@@
Appendix B Accounting fraud detection
@@
Figure 8 shows the prompt used to assess whether amendment reasons are related to accounting fraud violations. Figure 9 and 10 show examples of model responses.
Prompt to detect if the reason of the amendment is related to accounting fraud 以下のテキストは訂正有価証券報告書の冒頭部分です。この訂正有価証券報告書が不適切会計、粉飾決算、会計不正に関連しているかどうかを判断してください。 特に「提出理由」の部分に着目し、以下のような言葉や表現がある場合は不正会計の可能性が高いと考えられます: - 不適切会計 - 会計不正 - 不正行為 - 粉飾決算 - 会計処理の誤り - 売上の過大計上 - 費用の過少計上 - 資産の過大評価 - 不適切な収益認識 - 監査法人からの指摘 - 社内調査 - 第三者委員会 訂正の理由が単純な記載ミスや軽微な修正ではなく、会計上の重大な問題を示している場合は「Yes」と回答し、詳細な説明を提供してください。 特に財務諸表(貸借対照表、損益計算書、キャッシュフロー計算書など)の数値に変更が生じた事例に注目してください。 会計不正とは関係ない場合や財務諸表に変更が生じていない場合は「No」と回答し、その理由を簡潔に説明してください。 以下のJSON形式で回答してください。このJSONは必ず有効なJSON形式である必要があります: “‘json { ”is_accounting_fraud”: bool # true or false ”explanation”: str # 理由を説明する ”company_name”: str # 会社名 } “‘ 回答は必ずこの形式に一致させてください。 JSONは必ず上記の形式の“‘json“‘タグで囲んで提供してください。 訂正有価証券報告書のテキスト:
@@
Figure 9: Accounting fraud detection analysis with Claude 3.7 Sonnet. Input: BS+CF+PL+Summary+Text, Ground truth: non-fraud.
@@
Figure 10: Accounting fraud detection analysis with GPT-4o. Input: BS+CF+PL+Summary+Text, Ground truth: non-fraud.
Appendix C Earnings forecasting
Figure 11 and 12 shows examples of model response.
@@ Appendix C Earnings forecasting
Figure 11: Earnings forecasting analysis with Claude 3.7 Sonnet. Input: BS+CF+PL+Summary+Text, Ground truth: Increase.
@@
Figure 12: Earnings forecasting analysis with GPT-4o. Input: BS+CF+PL+Summary+Text, Ground truth: Increase.
Appendix D Industry prediction
Table 9 shows the list of the 33 and 17 class labels we use for industry prediction. Figure 13 and 14 show examples of model response.
Table 9: Mapping of SICC 33-sector classification to the TOPIX-17 Series.